# Bias and Fairness in Large Language Models: A Survey

### v2

@@ 4.1.3 Discussion and Limitations
Blodgett et al. (2021) highlight several severe shortcomings of several counterfactual input datasets, namely Winogender, WinoBias, StereoSet, and CrowS-Pairs. [^4] As a result of Blodgett et al. (2021), CrowS-Pairs now includes a data reliability disclaimer at https://github.com/nyu-mll/crows-pairs. In nearly half of all instances, these datasets contain ambiguities about what stereotypes they capture, raising questions whether they are valid indicators of real-world stereotypes at all. At the core of these issues are (i) unclear articulation of the power imbalances, harm, or other stereotype content present each instance; and (ii) inconsistent, invalid, or unrelated perturbations of social groups. For example, it is unclear how racial bias against Ethiopia is captured by StereoSet’s stereotype, anti-stereotype pair:
@@
Beyond data reliability, these datasets may also have limited generalizability to broader populations. These datasets are often situated in the United States context – for instance, occupation-gender datasets like Winogender, WinoBias, WinoBias+, and BEC-Pro leverage data from the U.S. Department of Labor – yet are offered as a general benchmark for English language everywhere. Datasets constructed by instantiating templates with protected attribute or other words may also lack diversity and may be unrepresentative of real use cases of the models.
@@ 4.2 Prompts
Though some sentence pair datasets can be adapted as input sentences to condition text continuation, other datasets have been proposed as prompts that specify the first few words in a sentence, or propose a question, and ask the model to provide a continuation or answer.
@@ 4.2.1 Sentence Completions
Instead of using contrived bias triggers, as often done in counterfactual input datasets, these datasets aim to match the distribution of human-written text to model more natural language use and potentially untargeted biases. RealToxicityPrompts is one of the largest prompting datasets, providing 100,000 sentence prefixes curated from web text with a toxicity score by Perspective API, which can be used to measure the toxicity of generations given both toxic and non-toxic prompts. To create the dataset, a set of web-scraped sentences are scored for toxicity, and 25K sentences are sampled from each of four quartiles, then split into a prompt (used in the dataset) and a continuation. BOLD introduces 23,679 prompts to assess bias in profession, gender, race, religion, and political ideology. The dataset is collected by scraping English Wikipedia pages that mention a group in the bias domain (i.e., profession), and truncating sentences to form prompts. HONEST (Nozza et al., 2021) provides 420 sentences to measure negative gender stereotypes in sentence completions in English, Italian, French, Portuguese, Spanish, and Romanian. Each sentence follows a cloze-based form, with a gendered identity term in the prompt, that be can completed by prompting a free-text continuation. Note that a completion may also be done by replacing a single [MASK] token, which enables compatibility with masked token-based metrics (Section 3.4.1), in addition to generated text-based ones. TrustGPT (Huang et al., 2023) provides prompts to evaluate toxicity and performance disparities between social groups. For toxicity evaluation, prompts ask the model to state something toxic, bad, or harmful, given some social norm; the completion can be scored for its level of toxicity. In contrast, for bias evaluation, prompts ask for a toxic completion given both a social group and social norm; completions can be compared between social groups to measure disparities in toxicity levels.
@@ 4.2.2 Question-Answering
With a related task, Gender Representation-Bias for Information Retrieval (Grep-BiasIR) (Krieg et al., 2023) provides 118 gender-neutral search queries for document retrieval to assess gender representation bias. Instead of providing associated answers as done with question-answering, Grep-BiasIR pairs each query with a relevant and non-relevant document with feminine, masculine, and neutral variations, with 708 documents in total. A disproportional retrieval of either feminine or masculine documents illustrates bias.
@@ 4.2.3 Discussion and Limitations
Akyürek et al. (2022) show that ambiguity may emerge when one social group is mentioned in a prompt, and another is mentioned in the completion, creating ambiguity about to whom the bias or harm should refer. In other words, this over-reliance on social group labels can create misleading or incomplete evaluations. Akyürek et al. (2022) suggests reframing prompts to introduce a situation, instead of a social group, and then examining the completion for social group identifiers. These datasets also suffer from some data reliability issues, but to a lesser extent than those discussed in Blodgett et al. (2021) (Bommasani et al., 2023).
@@ 5.1 Pre-Processing Mitigation
Pre-processing mitigations broadly encompass measures that affect model inputs – namely, data and prompts – and do not intrinsically change the model’s trainable parameters. These mitigations seek to create more representative training datasets by adding underrepresented examples to the data via data augmentation (Section 5.1.1), carefully curating or upweighting the most effective examples for debiasing via data filtering and reweighting (Section 5.1.2), generating new examples that meet a set of targeted criteria (Section 5.1.3), changing prompts fed to the model (Section 5.1.4), or debiasing pre-trained contextualized representations before fine-tuning (Section 5.1.5). A pre-trained model can be fine-tuned on the transformed data and prompts, or initialized with the transformed representations. Examples of these techniques are shown in Figure 7.
Figure 7: Example Pre-Processing Mitigation Techniques (§ 5.1). We provide examples of data augmentation, filtering, re-weighting, and generation on the left, as well as various types of instruction tuning on the right. The first example illustrates counterfactual data augmentation, flipping binary gender terms to their opposites. Data filtering illustrates the removal of biased instances, such as derogatory language. Reweighting demonstrates how instances representing underrepresented or minority instances may be upweighted for training. Data generation shows how new examples may be constructed by human or machine writers based on priming examples that illustrate the desired standards for the new data. Instruction tuning modifies the prompt fed to the model by appending addition tokens. In the first example of modified prompting language, positive triggers are added to the input to condition the model to generate more positive outputs (based on Abid et al. (2021) and Narayanan Venkit et al. (2023)). Control tokens in this example indicate the presence ($+$) or absence (00) of masculine $M$ or feminine $F$ characters in the sentence (based on Dinan et al. (2020)). Continuous prompt tuning prepends the prompt with trainable parameters $p_{1},\cdots,p_{m}$.
@@ 5.1.1 Data Augmentation
Interpolation.
Based on Zhang et al. (2018b)’s mixup technique, interpolation techniques interpolate counterfactually-augmented training examples with the original versions and their labels to extend the distribution of the training data. Ahn et al. (2022) leverage the mixup framework to equalize the pre-trained model’s output logits with respect to two opposing words in a gendered pair. Yu et al. (2023b) introduce Mix-Debias, and use mixup on an ensemble of corpora to reduce gender stereotypes with an expanded training set.
@@ 5.1.2 Data Filtering and Reweighting
Dataset filtering.
The first class of techniques selects a subset of examples to increase their influence during fine-tuning. Garimella et al. (2022) and Borchers et al. (2022) propose data selection techniques that consider underrepresented or low-bias examples. Garimella et al. (2022) curate and filter text written by historically disadvantaged gender, racial, and geographical groups for fine-tuning, to enable the model to learn more diverse world views and linguistic norms. Borchers et al. (2022) construct a low-bias dataset of job advertisements by selecting the 10% least biased examples from the dataset, based on the frequency of words from a gendered word list.
In contrast, other data selection methods focus on the most biased examples to neutralize or filter out. In a neutralizing approach for gender bias mitigation, Thakur et al. (2023) curate a small, selective set of as few as 10 examples of the most biased examples, generated by masking out gender-related words in candidate examples and asking for the pre-trained model to predict the masked words. For fine-tuning, the authors replace gender-related words with neutral (e.g., "they") or equalized (e.g., "he or she") alternatives.
Using instead a filtering approach, Raffel et al. (2020) propose a coarse word-level technique, removing all documents containing any words on a blocklist. Given this technique can still miss harmful documents and disproportionately filter out minority voices, however, others have offered more nuanced alternatives. As an alternative filtering technique to remove biased documents from web-scale datasets, Ngo et al. (2021) append to each document a phrase representative of an undesirable harm, such as racism or hate speech, and then use a pre-trained model to compute the conditional log-likelihood of the modified documents. Documents with high log-likelihoods are removed from the training set. Similarly, Sattigeri et al. (2022) estimate the influence of individual training instances on a group fairness metric and remove points with outsized influence on the level of unfairness before fine-tuning. Han et al. (2022a) downsamples majority-class instances to balance the number of examples in each class with respect to some protected attribute.
@@
Instance reweighting.
The second class of techniques reweights instances that should be (de)emphasized during training. Han et al. (2022a) use instance reweighting to equalize the weight of each class during training, calculating each instance’s weight in the loss as inversely proportional to its label and an associated protected attribute. Other approaches employed by Utama et al. (2020) and Orgad & Belinkov (2023) focus on downweighting examples containing social group information, even in the absence of explicit social group labels. Because bias factors are often surface-level characteristics that the pre-trained model uses as simple shortcuts for prediction, reducing the importance of stereotypical shortcuts may mitigate bias in fine-tuning. Utama et al. (2020) propose a self-debiasing method that uses a shallow model trained on a small subset of the data to identify potentially biased examples, which are subsequently downweighted by the main model during fine-tuning. Intuitively, the shallow model can capture similar stereotypical demographic-based shortcuts as the pre-trained model. Orgad & Belinkov (2023) also use an auxiliary classifier in their method BLIND to identify demographic-laden examples to downweight, but alternatively base the classifier on the predicted pre-trained model’s success.
Equalized teacher model probabilities.
Knowledge distillation is a training paradigm that transfers knowledge from a pre-trained teacher model to a smaller student model with fewer parameters. During distillation, the student model may inherit or even amplify biases from the teacher (Ahn et al., 2022; Silva et al., 2021). To mitigate this, the teacher’s predicted token probabilities can be modified before passing them to the student model as a pre-processing step. Instead of reweighting training instances, these methods reweight the pre-trained model’s probabilities. Delobelle & Berendt (2022) propose a set of user-specified probabilistic rules that can modify the teacher model’s outputs by equalizing the contextualized probabilities of two opposing gendered words given the same context. Gupta et al. (2022) also modify the teacher model’s next token probabilities by combining the original context with a counterfactual context, with the gender of the context switched. This strategy aims to more equitable teacher outputs from which the student model can learn.
@@ 5.1.3 Data Generation
Exemplary examples.
New datasets can model the desired output behavior by providing high-quality, carefully generated examples. Solaiman & Dennison (2021) present an iterative process to build a values-targeted dataset that reflects a set of topics (e.g., legally protected classes in the United States) from which to remove bias from the model. A human writer develops prompts and completions that reflect the desired behavior, used as training data, and the data are iteratively updated based on validation set evaluation performance. Also incorporating human writers, Dinan et al. (2020) investigate targeted data collection to reduce gender bias in chat dialogue models by curating human-written diversified examples, priming crowd workers with examples and standards for the desired data. Sun et al. (2023) construct example discussions that demonstrate and explain facets of morality, including fairness, using rules-of-thumb that encode moral principles and judgments. To train models that can appropriately respond and recover to biased input or outputs, Ung et al. (2022) generate a set of dialogues with example recovery statements, such as apologies, after unsafe, offensive, or inappropriate utterances. Similarly, Kim et al. (2022) generate a dataset of responses to biased or otherwise problematic statements based on rules-of-thumb that follow social norms.
Word lists.
Word-swapping techniques like CDA and CDS rely on word pair lists. Several works have presented word lists associated with social groups for gender (Bolukbasi et al., 2016; Garg et al., 2018; Gupta et al., 2022; Hall Maudslay et al., 2019; Lu et al., 2020; Zhao et al., 2017; 2018), race (Caliskan et al., 2017; Garg et al., 2018; Gupta et al., 2022; Manzini et al., 2019), age (Caliskan et al., 2017), dialect (Ziems et al., 2022), and other social group terms (Dixon et al., 2018). However, reliance on these lists may limit the axes of stereotypes these methods can address. To increase generality, Omrani et al. (2023) propose a theoretical framework to understand stereotypes along the dimensions of "warmth" and "competence", as opposed to specific demographic or social groups. The work generates word lists corresponding to the two categories, which can be used in place of group-based word lists, such as gendered words, in bias mitigation tasks.
@@ 5.1.4 Instruction Tuning
Modified prompting language.
Textual instructions or triggers may be added to a prompt to generate an unbiased output. Mattern et al. (2022) propose prompting language with different levels of abstraction to instruct the model to avoid using stereotypes. Narayanan Venkit et al. (2023) use adversarial triggers to mitigate nationality bias by prepending a positive adjective to the prompt to encourage more favorable perceptions of a country. This is similar to Abid et al. (2021), which prepend short phrases to prompt positive associations with Muslims to reduce anti-Muslim bias. Sheng et al. (2020) identify adversarial triggers that can induce positive biases for a given social group. The work iteratively searches over a set of input prompts that maximize neutral and positive sentiment towards a group, while minimizing negative sentiment.
Control tokens.
Instead of prepending instructive language to the input, control tokens corresponding to some categorization of the prompt can be added instead. Because the model learns to associate each control token with the class of inputs, the token can be set at inference to condition the generation. Dinan et al. (2020), for instance, mitigate gender bias in dialogue generation by binning each training example by the presence or absence of masculine or feminine gendered words, and appending a control token corresponding to the bin to each prompt. Xu et al. (2020) adapt Dinan et al. (2020)’s approach to provide safety guarantees for chatbot applications. The authors use safety control tokens, generated by a safety classifier that measures offensiveness, bias, and other potential harms in text. Similarly, Lu et al. (2022) score training examples with a reward function that quantifies some unwanted property, such as toxicity or bias, which is used to quantize the examples into bins. Corresponding reward tokens are prepended to the input.
Continuous prompt tuning.
Continuous prefix or prompt tuning (Lester et al., 2021; Li & Liang, 2021; Liu et al., 2021c) modifies the input with a trainable prefix. This technique freezes all original pre-trained model parameters and instead prepends additional trainable parameters to the input. Intuitively, the prepended tokens represent task-specific virtual tokens that can condition the generation of the output as before, but now enable scalable and tunable updates to task-specific requirements, rather than manual prompt engineering. As a bias mitigation technique, Fatemi et al. (2023) propose GEEP to use continuous prompt tuning to mitigate gender bias, fine-tuning on a gender-neutral dataset. In Yang et al. (2023)’s ADEPT technique, continuous prompts encourage neutral nouns and adjectives to be independent of protected attributes.

### v3

@@ 4.1.3 Discussion and Limitations
Blodgett et al. (2021) highlight several severe shortcomings of several counterfactual input datasets, namely Winogender, WinoBias, StereoSet, and CrowS-Pairs. [^4] As a result of Blodgett et al. (2021), CrowS-Pairs now includes a data reliability disclaimer at https://github.com/nyu-mll/crows-pairs. In nearly half of all instances, these datasets contain ambiguities about what stereotypes they capture, raising questions whether they are valid indicators of real-world stereotypes at all. At the core of these issues are (1) unclear articulation of the power imbalances, harm, or other stereotype content present in each instance; and (2) inconsistent, invalid, or unrelated perturbations of social groups. For example, it is unclear how racial bias against Ethiopia is captured by StereoSet’s stereotype, anti-stereotype pair:
@@
Beyond data reliability, these datasets may also have limited generalizability to broader populations. These datasets are often situated in the United States context — for instance, occupation-gender datasets like Winogender, WinoBias, WinoBias+, and BEC-Pro leverage data from the U.S. Department of Labor — yet are offered as a general benchmark for English language everywhere. Datasets constructed by instantiating templates with protected attribute or other words may also lack diversity and may be unrepresentative of real use cases of the models.
@@ 4.2 Prompts
Though some sentence pair datasets can be adapted as input sentences to condition text continuation, other datasets have been proposed as prompts that specify the start of a sentence, or propose a question, and ask the model to provide a continuation or answer. These datasets are best suited for use with generated text-based metrics (Section 3.5).
@@ 4.2.1 Sentence Completions
Instead of using contrived bias triggers, as often done in counterfactual input datasets, these datasets aim to match the distribution of human-written text to model more natural language use and potentially untargeted biases. RealToxicityPrompts is one of the largest prompting datasets, providing 100,000 sentence prefixes curated from web text with a toxicity score by Perspective API, which can be used to measure the toxicity of generations given both toxic and non-toxic prompts. To create the dataset, a set of web-scraped sentences are scored for toxicity, and 25K sentences are sampled from each of four quartiles, then split into a prompt (used in the dataset) and a continuation. BOLD introduces 23,679 prompts to assess bias in profession, gender, race, religion, and political ideology. The dataset is collected by scraping English Wikipedia pages that mention a group in the bias domain (i.e., profession), and truncating sentences to form prompts. HONEST (Nozza, Bianchi, and Hovy, 2021) provides 420 sentences to measure negative gender stereotypes in sentence completions in English, Italian, French, Portuguese, Spanish, and Romanian. Each sentence follows a cloze-based form, with a gendered identity term in the prompt, that can be completed by prompting a free-text continuation. Note that a completion may also be done by replacing a single [MASK] token, which enables compatibility with masked token-based metrics (Section 3.4.1), in addition to generated text-based ones. TrustGPT (Huang et al., 2023) provides prompts to evaluate toxicity and performance disparities between social groups. For toxicity evaluation, prompts ask the model to state something toxic or harmful, given some social norm; the completion can be scored for its level of toxicity. For disparity assessment, prompts ask for a toxic completion given both a social group and social norm; completions can be compared between social groups to measure disparities in toxicity levels.
@@ 4.2.2 Question-Answering
With a related task, Gender Representation-Bias for Information Retrieval (Grep-BiasIR) (Krieg et al., 2023) provides 118 gender-neutral search queries for document retrieval to assess gender representation bias. Instead of providing associated answers as done with question-answering, Grep-BiasIR pairs each query with a relevant and non-relevant document with feminine, masculine, and neutral variations, with 708 documents in total. A disproportional retrieval of feminine or masculine documents illustrates bias.
@@ 4.2.3 Discussion and Limitations
Akyürek et al. (2022) show that ambiguity may emerge when one social group is mentioned in a prompt, and another is mentioned in the completion, creating uncertainty about to whom the bias or harm should refer. In other words, this over-reliance on social group labels can create misleading or incomplete evaluations. Akyürek et al. (2022) suggests reframing prompts to introduce a situation, instead of a social group, and then examining the completion for social group identifiers. These datasets also suffer from some data reliability issues, but to a lesser extent than those discussed in Blodgett et al. (2021) (Liang et al., 2022).
4.3 Recommendations
We synthesize findings and guidance from the literature to make the following recommendations. For more detailed discussion and limitations, see Sections 4.1.3 and 4.2.3.
  1. 1.
    Exercise caution around construct, content, and ecological validity challenges. Rigorously assess whether the dataset clearly grounds and articulates the power imbalance it seeks to measure, and whether this articulation matches the targeted downstream bias. For datasets that rely on social group perturbations, verify that the counterfactual inputs accurately reflect real-world biases.
  2. 2.
    Ensure generalizability and applicability. Datasets should be selected to provide exhaustive coverage over a range of biases for multidimensional evaluation that extends beyond the most common axes of gender (identity) and stereotyping. Datasets constructed within specific contexts, such as the United States, should be used cautiously and limitedly as proxies for biases in other settings.
@@ 5.1 Pre-Processing Mitigation
Pre-processing mitigations broadly encompass measures that affect model inputs — namely, data and prompts — and do not intrinsically change the model’s trainable parameters. These mitigations seek to create more representative training datasets by adding underrepresented examples to the data via data augmentation (Section 5.1.1), carefully curating or upweighting the most effective examples for debiasing via data filtering and reweighting (Section 5.1.2), generating new examples that meet a set of targeted criteria (Section 5.1.3), changing prompts fed to the model (Section 5.1.4), or debiasing pre-trained contextualized representations before fine-tuning (Section 5.1.5). A pre-trained model can be fine-tuned on the transformed data and prompts, or initialized with the transformed representations. We show examples in Figure 7.
Figure 7: Example Pre-Processing Mitigation Techniques (§ 5.1). We provide examples of data augmentation, filtering, re-weighting, and generation on the left, as well as various types of instruction tuning on the right. The first example illustrates counterfactual data augmentation, flipping binary gender terms to their opposites. Data filtering illustrates the removal of biased instances, such as derogatory language (denoted as "@&!"). Reweighting demonstrates how instances representing underrepresented or minority instances may be upweighted for training. Data generation shows how new examples may be constructed by human or machine writers based on priming examples that illustrate the desired standards for the new data. Instruction tuning modifies the prompt fed to the model by appending additional tokens. In the first example of modified prompting language, positive triggers are added to the input to condition the model to generate more positive outputs (based on Abid, Farooqi, and Zou (2021) and Narayanan Venkit et al. (2023)). Control tokens in this example indicate the presence ($+$) or absence (00) of masculine $M$ or feminine $F$ characters in the sentence (based on Dinan et al. (2020)). Continuous prompt tuning prepends the prompt with trainable parameters $p_{1},\cdots,p_{m}$.
@@ 5.1.1 Data Augmentation
Interpolation
Based on Zhang et al. (2018)’s mixup technique, interpolation techniques interpolate counterfactually-augmented training examples with the original versions and their labels to extend the distribution of the training data. Ahn et al. (2022) leverage the mixup framework to equalize the pre-trained model’s output logits with respect to two opposing words in a gendered pair. Yu et al. (2023b) introduce Mix-Debias, and use mixup on an ensemble of corpora to reduce gender stereotypes.
@@ 5.1.2 Data Filtering and Reweighting
Dataset filtering
The first class of techniques selects a subset of examples to increase their influence during fine-tuning. Garimella, Mihalcea, and Amarnath (2022) and Borchers et al. (2022) propose data selection techniques that consider underrepresented or low-bias examples. Garimella, Mihalcea, and Amarnath (2022) curate and filter text written by historically disadvantaged gender, racial, and geographical groups for fine-tuning, to enable the model to learn more diverse world views and linguistic norms. Borchers et al. (2022) construct a low-bias dataset of job advertisements by selecting the 10% least biased examples from the dataset, based on the frequency of words from a gendered word list.
In contrast, other data selection methods focus on the most biased examples to neutralize or filter out. In a neutralizing approach for gender bias mitigation, Thakur et al. (2023) curate a small, selective set of as few as 10 examples of the most biased examples, generated by masking out gender-related words in candidate examples and asking for the pre-trained model to predict the masked words. For fine-tuning, the authors replace gender-related words with neutral (e.g., "they") or equalized (e.g., "he or she") alternatives. Using instead a filtering approach, Raffel et al. (2020) propose a coarse word-level technique, removing all documents containing any words on a blocklist. Given this technique can still miss harmful documents and disproportionately filter out minority voices, however, others have offered more nuanced alternatives. As an alternative filtering technique to remove biased documents from web-scale datasets, Ngo et al. (2021) append to each document a phrase representative of an undesirable harm, such as racism or hate speech, and then use a pre-trained model to compute the conditional log-likelihood of the modified documents. Documents with high log-likelihoods are removed from the training set. Similarly, Sattigeri et al. (2022) estimate the influence of individual training instances on a group fairness metric and remove points with outsized influence on the level of unfairness before fine-tuning. Han, Baldwin, and Cohn (2022a) downsamples majority-class instances to balance the number of examples in each class with respect to some protected attribute.
@@
Instance reweighting
The second class of techniques reweights instances that should be (de)emphasized during training. Han, Baldwin, and Cohn (2022a) use instance reweighting to equalize the weight of each class during training, calculating each instance’s weight in the loss as inversely proportional to its label and an associated protected attribute. Other approaches employed by Utama, Moosavi, and Gurevych (2020) and Orgad and Belinkov (2023) focus on downweighting examples containing social group information, even in the absence of explicit social group labels. Because bias factors are often surface-level characteristics that the pre-trained model uses as simple shortcuts for prediction, reducing the importance of stereotypical shortcuts may mitigate bias in fine-tuning. Utama, Moosavi, and Gurevych (2020) propose a self-debiasing method that uses a shallow model trained on a small subset of the data to identify potentially biased examples, which are subsequently downweighted by the main model during fine-tuning. Intuitively, the shallow model can capture similar stereotypical demographic-based shortcuts as the pre-trained model. Orgad and Belinkov (2023) also use an auxiliary classifier in their method BLIND to identify demographic-laden examples to downweight, but alternatively base the classifier on the predicted pre-trained model’s success.
Equalized teacher model probabilities
Knowledge distillation is a training paradigm that transfers knowledge from a pre-trained teacher model to a smaller student model with fewer parameters. In contrast to data augmentation, which applies to a fixed training dataset, knowledge distillation applies to the outputs of the teacher model, which may be dynamic in nature and encode implicit behaviors already learned by the model. During distillation, the student model may inherit or even amplify biases from the teacher (Ahn et al., 2022; Silva, Tambwekar, and Gombolay, 2021). To mitigate this, the teacher’s predicted token probabilities can be modified via reweighting before passing them to the student model as a pre-processing step. Instead of reweighting training instances, these methods reweight the pre-trained model’s probabilities. Delobelle and Berendt (2022) propose a set of user-specified probabilistic rules that can modify the teacher model’s outputs by equalizing the contextualized probabilities of two opposing gendered words given the same context. Gupta et al. (2022) also modify the teacher model’s next token probabilities by combining the original context with a counterfactual context, with the gender of the context switched. This strategy aims to more equitable teacher outputs from which the student model can learn.
@@ 5.1.3 Data Generation
Exemplary examples
New datasets can model the desired output behavior by providing high-quality, carefully generated examples. Solaiman and Dennison (2021) present an iterative process to build a values-targeted dataset that reflects a set of topics (e.g., legally protected classes in the United States) from which to remove bias from the model. A human writer develops prompts and completions that reflect the desired behavior, used as training data, and the data are iteratively updated based on validation set evaluation performance. Also incorporating human writers, Dinan et al. (2020) investigate targeted data collection to reduce gender bias in chat dialogue models by curating human-written diversified examples, priming crowd workers with examples and standards for the desired data. Sun et al. (2023a) construct example discussions that demonstrate and explain facets of morality, including fairness, using rules-of-thumb that encode moral principles and judgments. To train models that can appropriately respond and recover to biased input or outputs, Ung, Xu, and Boureau (2022) generate a set of dialogues with example recovery statements, such as apologies, after unsafe, offensive, or inappropriate utterances. Similarly, Kim et al. (2022) generate a dataset of prosocial responses to biased or otherwise problematic statements based on crowdsourced rules-of-thumb from the Social Chemistry dataset (Forbes et al., 2020) that represent socio-normative judgments.
Word lists
Word-swapping techniques like CDA and CDS rely on word pair lists. Several works have presented word lists associated with social groups for gender (Bolukbasi et al., 2016; Garg et al., 2018; Gupta et al., 2022; Hall Maudslay et al., 2019; Lu et al., 2020; Zhao et al., 2017, 2018), race (Caliskan, Bryson, and Narayanan, 2017; Garg et al., 2018; Gupta et al., 2022; Manzini et al., 2019), age (Caliskan, Bryson, and Narayanan, 2017), dialect (Ziems et al., 2022), and other social group terms (Dixon et al., 2018). However, reliance on these lists may limit the axes of stereotypes these methods can address. To increase generality, Omrani et al. (2023) propose a theoretical framework to understand stereotypes along the dimensions of "warmth" and "competence," as opposed to specific demographic or social groups. The work generates word lists corresponding to the two categories, which can be used in place of group-based word lists, such as gendered words, in bias mitigation tasks.
@@ 5.1.4 Instruction Tuning
Modified prompting language
Textual instructions or triggers may be added to a prompt to generate an unbiased output. Mattern et al. (2022) propose prompting language with different levels of abstraction to instruct the model to avoid using stereotypes. Similar to counterfactual augmentation, but distinct in their more generic application at the prompting level (as opposed to specific perturbations for each data instance), Narayanan Venkit et al. (2023) use adversarial triggers to mitigate nationality bias by prepending a positive adjective to the prompt to encourage more favorable perceptions of a country. This is similar to Abid, Farooqi, and Zou (2021), which prepend short phrases to prompt positive associations with Muslims to reduce anti-Muslim bias. Sheng et al. (2020) identify adversarial triggers that can induce positive biases for a given social group. The work iteratively searches over a set of input prompts that maximize neutral and positive sentiment towards a group, while minimizing negative sentiment.
Control tokens
Instead of prepending instructive language to the input, control tokens corresponding to some categorization of the prompt can be added instead. Because the model learns to associate each control token with the class of inputs, the token can be set at inference to condition the generation. Dinan et al. (2020), for instance, mitigate gender bias in dialogue generation by binning each training example by the presence or absence of masculine or feminine gendered words, and appending a control token corresponding to the bin to each prompt. Xu et al. (2020) adapt this approach to reduce offensive language in chatbot applications. The authors identify control tokens using a classifier that measures offensiveness, bias, and other potential harms in text. The control tokens can be appended to the input during inference to control model generation. Similarly, Lu et al. (2022) score training examples with a reward function that quantifies some unwanted property, such as toxicity or bias, which is used to quantize the examples into bins. Corresponding reward tokens are prepended to the input.
Continuous prompt tuning
Continuous prefix or prompt tuning (Lester, Al-Rfou, and Constant, 2021; Li and Liang, 2021; Liu et al., 2021c) modifies the input with a trainable prefix. This technique freezes all original pre-trained model parameters and instead prepends additional trainable parameters to the input. Intuitively, the prepended tokens represent task-specific virtual tokens that can condition the generation of the output as before, but now enable scalable and tunable updates to task-specific requirements, rather than manual prompt engineering. As a bias mitigation technique, Fatemi et al. (2023) propose GEEP to use continuous prompt tuning to mitigate gender bias, fine-tuning on a gender-neutral dataset. In Yang et al. (2023)’s ADEPT technique, continuous prompts encourage neutral nouns and adjectives to be independent of protected attributes.